5(2x-7)+3(4x-7)-(x+2)=-142
I need help on how to solve this problem??

Answer by Edwin McCravy(20056)


5(2x-7) + 3(4x-7) - (x+2) = -142

Put a 1 before the (x+2)

5(2x-7) + 3(4x-7) - 1(x+2) = -142

Distribute:

10x - 35 + 12x - 21 - 1x - 2 = -142

Combine 10x + 12x - 1x to get 21x
Combine -35 - 21 - 2 to get -58

                    21x - 58 = -142

Add 58 to both sides:

                         21x = -84

Divide both side by 21

                    21x/21 = -84/21 
             
                         x = -4

Edwin